### v2

# TSQCA: Threshold-Sweep Qualitative Comparative Analysis in R

Abstract
Qualitative Comparative Analysis (QCA) requires researchers to choose calibration and dichotomization thresholds, and these choices can substantially affect truth tables, minimization, and resulting solution formulas. Despite this dependency, threshold sensitivity is often examined only in an ad hoc manner because repeated analyses are time-intensive and error-prone. We present TSQCA, an R package that automates threshold-sweep analyses by treating thresholds as explicit analytical variables. It provides four sweep functions (otSweep, ctSweepS, ctSweepM, dtSweep) to explore outcome thresholds, single-condition thresholds, multi-condition threshold grids, and joint outcome–condition threshold spaces, respectively. TSQCA integrates with the established CRAN package QCA for truth table construction and Boolean minimization, while returning structured S3 objects with consistent print/summary methods and optional detailed results. The package also supports automated Markdown report generation and configuration-chart output to facilitate reproducible documentation of cross-threshold results.
Keywords: Qualitative Comparative Analysis, QCA, Threshold sensitivity, R package, Configurational analysis
1 Introduction
1.1 The threshold dependency problem in QCA
Qualitative Comparative Analysis (QCA) has become a standard method for analyzing complex causal relationships in social science research (Ragin, 1987, 2008; Schneider and Wagemann, 2012). QCA can handle equifinality and conjunctural causation by identifying combinations of conditions that are sufficient (or necessary) for an outcome.
A fundamental and critical step in QCA is calibrating raw data into set membership scores. In crisp-set QCA (csQCA), this process involves dichotomizing continuous variables using thresholds. Even in fuzzy-set QCA (fsQCA), calibration anchors (full membership, crossover point, and full non-membership) similarly require threshold decisions (Schneider and Wagemann, 2012).
Despite the crucial role of thresholds, systematic examination of their impact remains limited. Researchers typically report results for a single threshold setting, and robustness checks—when conducted—are often relegated to appendices. This practice obscures a critical question: How stable are QCA findings to reasonable threshold variations? Previous research has demonstrated that QCA results can be highly sensitive to parameter choices, including calibration thresholds (Krogslund et al., 2015; Thiem et al., 2016).
This issue is not merely technical. In applied research, thresholds often reflect substantive decisions about what constitutes “high” or “low” values. Whether the threshold for “high customer satisfaction” is set at 7 or 8 on a satisfaction scale carries different implications. TSQCA enables researchers to examine how such decisions affect their conclusions.
1.2 Existing approaches and their limitations
Current QCA software provides limited functionality for threshold sensitivity analysis. Table 1 presents a comparison of major packages.
Table 1: Threshold sensitivity features in existing packages
Package Threshold Sensitivity Feature Limitation
QCA (Duşa, 2019) Manual recalibration required Systematic exploration difficult
SetMethods (Oana and Schneider, 2018) Sensitivity range for calibration anchors Binary thresholds not covered
fs/QCA (Ragin and Davey, 2016) No feature
Tosmana (Cronqvist, 2019) No feature
The SetMethods package provides functionality to assess sensitivity ranges for consistency cutoffs and calibration anchors following the robustness test protocol proposed by Oana and Schneider. These tools primarily focus on evaluating local robustness around a given calibration or cutoff choice. However, no existing package offers a systematic sweep-based exploration of dichotomization thresholds in csQCA across a predefined threshold space. As a result, researchers typically rely on manually repeating analyses for different threshold settings, a process that is both time-consuming and prone to error.
1.3 Contributions of TSQCA
The TSQCA package addresses this methodological gap in six ways:
  1. 1.
    Automation of threshold sweeps: A single function call explores the entire threshold range specified by the user. This enables efficient execution of analyses across a large, predefined grid of threshold settings.
  2. 2.
    Structured output: Results are organized as S3 objects with print() and summary() methods, facilitating cross-threshold comparisons. Summary tables, detailed solution objects, and reproducibility settings are included.
  3. 3.
    Integration with the QCA package: TSQCA uses the core functions of the established QCA package (Duşa, 2019) internally, including support for negated outcomes (outcome = "~Y"). This ensures methodological consistency and allows researchers to benefit from updates to the QCA package.
  4. 4.
    Report generation: Structured markdown reports are automatically generated from analysis results, improving reproducibility and ease of reporting.
  5. 5.
    Configuration table generation: Fiss-style configuration tables (Fiss, 2011) with standard QCA notation ($\bullet$ for presence, $\otimes$ for absence) can be generated automatically, supporting Unicode, ASCII, and LaTeX output formats for direct use in academic publications.
  6. 6.
    Multiple solution handling: For multiple equivalent solutions, the package identifies Essential Prime Implicants (EPIs) and Selective Prime Implicants (SPIs).
The package transforms thresholds from fixed parameters to analytical variables, enabling researchers to systematically map the solution space. This shift in perspective—from treating thresholds as “assumptions” to treating them as “objects of study”—opens new analytical possibilities.
1.4 Paper organization
The remainder of this paper is organized as follows. Section 2 reviews the set-theoretic foundations of QCA and clarifies why calibration thresholds constitute a critical source of analytical dependency. Section 3 describes the design philosophy of the TSQCA package and its relationship to the established QCA package. Section 4 introduces the main functions of TSQCA and demonstrates their basic usage through fully reproducible examples based on the bundled sample dataset. Section 5 presents the advanced output structures and reporting capabilities of TSQCA, illustrating how results from threshold-sweep analyses can be systematically summarized, visualized, and exported for reproducible reporting. Section 6 compares TSQCA with existing packages. Section 7 discusses scope, limitations, and potential extensions. Section 8 concludes the paper.
2 Theoretical background
2.1 Calibration and threshold dependency
This section does not aim to provide a comprehensive introduction to QCA. Instead, it focuses exclusively on the set-theoretic concepts necessary to understand threshold sweep analysis as implemented in TSQCA. Readers seeking a full theoretical treatment of QCA are referred to established textbooks and reviews (Schneider and Wagemann, 2012; Oana et al., 2021; Thomann and Maggetti, 2020).
QCA examines subset relationships between condition configurations and outcomes, quantified through consistency and coverage indices (Ragin, 2006; Schneider and Wagemann, 2012). For a comprehensive treatment of set-theoretic foundations and their application in QCA, see Schneider and Wagemann (2012) and Duşa (2019).
A critical step in crisp-set QCA is dichotomizing continuous variables using thresholds:
[…]
Stage 1: Set membership determination
[…]
Stage 2: Truth table construction
Changes in set membership alter the number of cases (n) and the consistency score (incl) for each configuration in the truth table. Specifically:
  • Configurations falling below the frequency threshold (n.cut) are excluded from analysis
  • Configurations falling below the consistency threshold (incl.cut) are treated as outcome=0
Stage 3: Boolean minimization
Changes to the truth table yield different prime implicants through the Quine-McCluskey algorithm or enhanced methods such as Consistency Cubes (Duşa, 2018), altering the final solution expression.
2.2 The concept of hierarchical sufficiency
When varying outcome thresholds, we often observe a pattern we call “hierarchical sufficiency”: as the required outcome level increases, the configurations sufficient for that outcome become more complex, while their number decreases.
Figure 1 illustrates this concept schematically. At lower thresholds (thrY = 5), multiple relatively simple configurations may suffice—indicated by multiple paths (M1, M2, M3). As thresholds increase, paths begin to disappear, and remaining paths require more conditions. At very high thresholds (thrY = 9), no configuration may meet the consistency requirements.
Figure 1: Conceptual illustration of hierarchical sufficiency across outcome thresholds. As the outcome threshold increases from thrY = 5 to thrY = 9, the number of sufficient paths decreases from three (M1, M2, M3) to none. This figure is schematic and intended to illustrate the general logic of hierarchical sufficiency; actual empirical results vary by dataset.
This pattern aligns with the intuitive understanding that achieving higher outcome levels requires more demanding combinations of conditions. Hierarchical sufficiency is also practically important, because it yields strategic implications according to target levels. For moderate targets, simple interventions may suffice, whereas ambitious targets typically require more comprehensive approaches.
We introduce this concept as a descriptive pattern observed through threshold sweep analysis, not as a formal theoretical law.
2.3 Relationship to existing robustness research
Discussion of robustness in QCA has been developed most prominently by Skaaning (2011) and Oana and Schneider (2021). Related work on sensitivity diagnostics have sought to formalize how QCA results respond to parameter perturbations (Thiem et al., 2016). More recently, Baumgartner (2022) investigated the concept of robust sufficiency and, through simulation experiments, benchmarked the performance of different QCA solution types in recovering robustly sufficient conditions.
While Baumgartner (2022) evaluates whether different QCA solution types can reliably recover robustly sufficient conditions, TSQCA does not seek to establish robustness in this sense; instead, it offers a systematic and descriptive mapping of how sufficient configurations vary across substantively meaningful outcome and condition thresholds.
Skaaning (2011) examined how variations in frequency and consistency thresholds affect QCA results and highlighted the importance of assessing solution stability.
Oana and Schneider (2021) proposed a more comprehensive robustness test protocol, implemented in the SetMethods package. This protocol includes:
  • Sensitivity range: the range over which parameters can be varied without altering the solution structure
  • Fit-oriented robustness: Variation in consistency and coverage
  • Case-oriented robustness: stability of case membership and assignment across solutions
However, these existing approaches have a confirmatory character. That is, they evaluate how robust results are to the thresholds set by the researcher.
In contrast, TSQCA adopts an exploratory approach. It treats thresholds as “manipulable variables” rather than “fixed values,” revealing how causal structures are generated, disappear, branch, or reorganize as thresholds change. Table 2 illustrates this difference.
Importantly, TSQCA is not designed as a robustness testing tool in the confirmatory sense. Rather than evaluating whether a single solution remains invariant within a tolerance range, we adopt an exploratory perspective in which thresholds are treated as manipulable analytical variables. This perspective allows researchers to examine how sufficient condition structures emerge, transform, and disappear across substantively meaningful threshold levels.
Table 2: Comparison of confirmatory and exploratory approaches
Aspect Confirmatory (Existing) Exploratory (TSQCA)
Meaning of threshold Range of plausible calibration choices Qualitatively different theoretical levels
Interpretation of change Problem (vulnerability) Analytical opportunity
Threshold selection Single theoretically justified threshold Compare across multiple thresholds
Purpose Validate initial solution Discover threshold-dependent causal paths
3 Package design
3.1 Design philosophy
TSQCA is designed according to four fundamental principles:
Principle 1: Complementarity
TSQCA builds upon, rather than replaces, existing QCA packages. It uses the truthTable() and minimize() functions provided by the QCA package (Thiem and Duşa, 2013; Duşa, 2019) directly, adding a new layer of threshold exploration. This design ensures:
  • Consistency with QCA methodology is ensured
  • Automatic benefits from QCA package updates (algorithm improvements, etc.)
  • Output format is familiar to QCA users
Principle 2: Transparency
All intermediate results are made available, enabling verification of the analysis process. Truth tables, solution objects, necessity analysis results at each threshold are saved and accessible.
Principle 3: Reproducibility
Complete reconstruction of analyses is enabled. All setting parameters are saved in result objects, and generate_report() automatically generates reports that include reproducibility information.
Principle 4: Task-specific interfaces
Although a single general-purpose function could technically handle all sweep types through parameter configuration, TSQCA provides dedicated functions for each analytical task: outcome threshold sweeps, single-condition sweeps, multi-condition sweeps, and dual sweeps. This design reduces cognitive load, improves computational efficiency by avoiding unnecessary grid expansion, and ensures that output formats are optimized for each specific use case.
From a user perspective, this design minimizes the cognitive and technical cost of conducting large-scale threshold sensitivity analyses, reducing what would otherwise require dozens of manual re-analyses to a single, reproducible function call.
3.2 Three types of QCA solutions
QCA minimization can produce three types of solutions depending on how logical remainders (unobserved configurations) are handled. Following standard QCA terminology (Schneider and Wagemann, 2012; Duşa, 2019), TSQCA supports all three solution types through the include and dir.exp parameters:
Table 3: Three types of QCA solutions and their parameter settings
Solution Type include dir.exp Description
Complex (Conservative) "" NULL Most conservative; uses only observed configurations
Parsimonious "?" NULL Most simplified; uses all logical remainders
Intermediate "?" c(1,1,...) Theory-guided; uses remainders consistent with directional expectations
As of version 1.1.0, TSQCA uses the same default parameter values as QCA::minimize(): include = "" and dir.exp = NULL. This produces the complex (conservative) solution by default, ensuring consistency with the established QCA package.
Researchers can obtain other solution types by explicitly specifying these parameters:
# Complex solution (default)
result <- otSweep(dat, outcome, conditions, sweep_range, thrX)
# Parsimonious solution
result <- otSweep(dat, outcome, conditions, sweep_range, thrX,
                  include = "?")
# Intermediate solution
result <- otSweep(dat, outcome, conditions, sweep_range, thrX,
                  include = "?", dir.exp = c(1, 1, 1))
3.3 Relationship to the QCA package
Figure 2 illustrates the TSQCA workflow and its integration with the QCA package. TSQCA operates as a wrapper that iteratively calls QCA package functions for each threshold value and aggregates results into a structured S3 object.
Figure 2: TSQCA workflow diagram. The package takes raw data and threshold range specifications as input, then iterates through each threshold value. At each iteration, TSQCA calls QCA::truthTable() and QCA::minimize() from the QCA package to perform standard QCA analysis. Results are aggregated across all thresholds and returned as an S3 object containing both a summary table for cross-threshold comparison and detailed results for each threshold.
Below is simplified R code illustrating this relationship:
otSweep <- function(dat, outcome, conditions, sweep_range, thrX,
                    incl.cut = 0.8, n.cut = 1, ...) {
  results <- list()
  for (t in sweep_range) {
    # Step 1: Dichotomize data at threshold t
    dat_binary <- dichotomize_data(dat, outcome, t, conditions, thrX)
    # Step 2: Call QCA package functions
    tt <- QCA::truthTable(
      data = dat_binary,
      outcome = outcome,
      conditions = conditions,
      incl.cut = incl.cut,
      n.cut = n.cut,
    sol <- QCA::minimize(tt, ...)
    # Step 3: Extract and save results
    results[[as.character(t)]] <- list(
      threshold = t,
      truth_table = tt,
      solution = sol,
      metrics = extract_metrics(sol)
  # Build summary table
  summary_table <- build_summary(results)
  structure(
    list(
      summary = summary_table,
      details = results,
      settings = list(sweep_range = sweep_range, thrX = thrX, ...)
    class = "otSweep"
Through this design, TSQCA functions as a “meta-analysis framework that repeatedly calls QCA.”
3.4 Package structure
The TSQCA package has the following structure:
TSQCA/
|-- R/
|   |-- otSweep.R         # Outcome threshold sweep
|   |-- ctSweepS.R        # Single condition threshold sweep
|   |-- ctSweepM.R        # Multiple condition threshold sweep
|   |-- dtSweep.R         # Dual threshold sweep
|   |-- generate_report.R # Report generation
|   |-- utils.R           # Utility functions
|   ‘-- data.R            # Dataset documentation
|-- data/
|   ‘-- sample_data.rda   # Sample data
|-- man/                  # Roxygen2-generated documentation
|-- vignettes/
|   |-- TSQCA_Tutorial_EN.Rmd
|   ‘-- TSQCA_Tutorial_JA.Rmd
|-- tests/
|   ‘-- testthat/         # Unit tests
|-- DESCRIPTION
|-- NAMESPACE
‘-- README.md
3.5 S3 methods
The use of S3 classes allows users to treat sweep results as analyzable objects, enabling downstream comparison, visualization, and reporting beyond printed solutions.
TSQCA implements S3 methods for all result objects:
### print method: Display summary table
print.otSweep <- function(x, ...) {
[…]
  cat("Sweep range:", min(x\$settings\$sweep_range), "-",
[…]
  cat("Condition thresholds:",
      paste(names(x\$settings\$thrX), "=", x\$settings\$thrX,
[…]
  print(x\$summary)
  invisible(x)
### summary method: Detailed statistical information
summary.otSweep <- function(object, ...) {
[…]
  # Solution stability analysis
  n_thresholds <- nrow(object\$summary)
  unique_solutions <- length(unique(object\$summary\$expression))
[…]
  cat("Solution stability:",
[…]
  # Consistency and coverage ranges
  cat("Consistency range:",
[…]
  cat("Coverage range:",
[…]
  invisible(object)
3.6 Function overview
Table 4 summarizes the main functions provided by TSQCA.
Table 4: TSQCA function overview
Function Purpose Key Arguments
otSweep() Outcome threshold sweep sweep_range, thrX, include, dir.exp
ctSweepS() Single condition sweep sweep_var, sweep_range, thrY, dir.exp
ctSweepM() Multi-condition sweep sweep_list, thrY, include, dir.exp
dtSweep() Dual threshold sweep sweep_range_Y, sweep_list_X, dir.exp
generate_report() Report generation result, output_file
3.7 Dependencies and imports
TSQCA has minimal dependencies to ensure stability:
Imports:
    QCA (>= 3.0),
    stats,
    utils
Suggests:
    testthat (>= 3.0.0),
    knitr,
    rmarkdown
The package imports only the QCA package as a hard dependency, ensuring that users can install and use TSQCA with minimal additional packages.
4 Main functions
This section introduces the main functions provided by the TSQCA package. All examples use the dataset bundled with the package (sample_data) and are designed as minimal, fully reproducible demonstrations. Default consistency and frequency cutoffs are used unless explicitly specified. Detailed outputs are computed internally but omitted from display for brevity.
Note: The examples in this section compute the intermediate solution by explicitly specifying include = "?" and dir.exp = c(1, 1, 1). To obtain the complex solution (default in v1.1.0), omit these parameters.
library(TSQCA)
data("sample_data", package = "TSQCA")
4.1 Outcome Threshold Sweep: otSweep()
4.1.1 Purpose
The otSweep() function explores how sufficient condition structures change as the outcome threshold varies, while condition thresholds are held constant. This function typically serves as the starting point for threshold-sweep analysis.
4.1.2 Function signature
otSweep(
  dat,
  outcome,
  conditions,
  sweep_range,
  thrX,
  incl.cut = 0.8,
  n.cut = 1,
  include = "",
  dir.exp = NULL,
  return_details = FALSE
4.1.3 Example with bundled sample data
res_ots <- otSweep(
  dat            = sample_data,
  outcome        = "Y",
  conditions     = c("X1", "X2", "X3"),
  sweep_range    = 6:8,
  thrX           = c(X1 = 7, X2 = 7, X3 = 7),
  include        = "?",
  dir.exp        = c(1, 1, 1),
  return_details = TRUE
summary(res_ots)
#> OTS-QCA Summary
#> Analysis Parameters:
#>   Outcome: Y
#>   Conditions: X1, X2, X3
#>   Consistency cutoff: 0.8
#>   Frequency cutoff: 1
#> Results by Threshold:
#>  thrY  expression inclS  covS n_solutions
#>     6  X3 + X1*X2 0.906 0.853           1
#>     7  X3 + X1*X2 0.906 0.879           1
#>     8 No solution    NA    NA           0
# Detailed results are computed internally
# but omitted from the output for brevity.
4.2 Single-Condition Threshold Sweep: ctSweepS()
4.2.1 Purpose
The ctSweepS() function examines sensitivity to the threshold of a single condition, while holding all other condition thresholds and the outcome threshold fixed.
4.2.2 Function signature
ctSweepS(
  dat,
  outcome,
  conditions,
  thrY,
  sweep_var,
  sweep_range,
  thrX_default,
  incl.cut = 0.8,
  n.cut = 1,
  include = "",
  dir.exp = NULL,
  return_details = FALSE
4.2.3 Example
res_cts <- ctSweepS(
  dat            = sample_data,
  outcome        = "Y",
  conditions     = c("X1", "X2", "X3"),
  sweep_var      = "X3",
  sweep_range    = 6:9,
  thrY           = 7,
  thrX_default   = 7,
  include        = "?",
  dir.exp        = c(1, 1, 1),
  return_details = TRUE
summary(res_cts)
#> CTS-QCA Summary
#> Analysis Parameters:
#>   Outcome: Y
#>   Conditions: X1, X2, X3
#>   Consistency cutoff: 0.8
#>   Frequency cutoff: 1
#> Results by Threshold:
#>  threshold expression inclS  covS n_solutions
#>          6      X1*X2 1.000 0.303           1
#>          7 X3 + X1*X2 0.906 0.879           1
#>          8 X3 + X1*X2 1.000 0.818           1
#>          9 X3 + X1*X2 1.000 0.515           1
4.3 Multiple Condition Threshold Sweep: ctSweepM()
4.3.1 Purpose
The ctSweepM() function simultaneously varies thresholds for multiple conditions, enabling exploration of the joint threshold space.
4.3.2 Function signature
ctSweepM(
  dat,
  outcome,
  conditions,
  thrY,
  sweep_list,
  incl.cut = 0.8,
  n.cut = 1,
  include = "",
  dir.exp = NULL,
  return_details = FALSE
4.3.3 Example
res_mcts <- ctSweepM(
  dat            = sample_data,
  outcome        = "Y",
  conditions     = c("X1", "X2", "X3"),
  sweep_list     = list(
    X1 = 6:7,
    X2 = 6:7,
    X3 = 6:7
  thrY           = 7,
  include        = "?",
  dir.exp        = c(1, 1, 1),
  return_details = TRUE
summary(res_mcts)
#> MCTS-QCA Summary
#> Analysis Parameters:
#>   Outcome: Y
#>   Conditions: X1, X2, X3
#>   Consistency cutoff: 0.8
#>   Frequency cutoff: 1
#> Results by Threshold:
#>         threshold combo_id     expression inclS  covS n_solutions
#>  X1=6, X2=6, X3=6        1          X1*X2 0.833 0.455           1
#>  X1=7, X2=6, X3=6        2          X1*X2 1.000 0.394           1
#>  X1=6, X2=7, X3=6        3      X1*X2*~X3 0.818 0.273           1
#>  X1=7, X2=7, X3=6        4          X1*X2 1.000 0.303           1
#>  X1=6, X2=6, X3=7        5             X3 0.864 0.576           1
#>  X1=7, X2=6, X3=7        6 ~X1*X3 + X1*X2 0.931 0.818           1
#>  X1=6, X2=7, X3=7        7             X3 0.864 0.576           1
#>  X1=7, X2=7, X3=7        8     X3 + X1*X2 0.906 0.879           1
4.4 Dual Threshold Sweep: dtSweep()
4.4.1 Purpose
The dtSweep() function jointly varies outcome thresholds and condition thresholds, enabling two-dimensional exploration of the threshold space.
As the most general sweep function, dtSweep() can technically replicate the behavior of the other sweep functions through appropriate parameter settings. However, as discussed in Principle 4, we provide dedicated functions for simpler tasks in order to reduce cognitive load and improve computational efficiency.
4.4.2 Function signature
dtSweep(
  dat,
  outcome,
  conditions,
  sweep_list_X,
  sweep_range_Y,
  incl.cut = 0.8,
  n.cut = 1,
  include = "",
  dir.exp = NULL,
  return_details = FALSE
4.4.3 Example
res_dts <- dtSweep(
  dat            = sample_data,
  outcome        = "Y",
  conditions     = c("X1", "X2", "X3"),
  sweep_list_X   = list(
    X1 = 6:7,
    X2 = 6:7,
    X3 = 6:7
  sweep_range_Y  = 6:7,
  include        = "?",
  dir.exp        = c(1, 1, 1),
  return_details = TRUE
summary(res_dts)
#> DTS-QCA Summary
#> Analysis Parameters:
#>   Outcome: Y
#>   Conditions: X1, X2, X3
#>   Consistency cutoff: 0.8
#>   Frequency cutoff: 1
#> Results by Threshold:
#>  thrY combo_id             thrX     expression inclS  covS n_solutions
#>     6        1 X1=6, X2=6, X3=6          X1*X2 0.833 0.441           1
#>     7        1 X1=6, X2=6, X3=6          X1*X2 0.833 0.455           1
#>     6        2 X1=7, X2=6, X3=6          X1*X2 1.000 0.382           1
#>     7        2 X1=7, X2=6, X3=6          X1*X2 1.000 0.394           1
#>     6        3 X1=6, X2=7, X3=6      X1*X2*~X3 0.818 0.265           1
#>     7        3 X1=6, X2=7, X3=6      X1*X2*~X3 0.818 0.273           1
#>     6        4 X1=7, X2=7, X3=6          X1*X2 1.000 0.294           1
#>     7        4 X1=7, X2=7, X3=6          X1*X2 1.000 0.303           1
#>     6        5 X1=6, X2=6, X3=7             X3 0.864 0.559           1
#>     7        5 X1=6, X2=6, X3=7             X3 0.864 0.576           1
#>     6        6 X1=7, X2=6, X3=7 ~X1*X3 + X1*X2 0.931 0.794           1
#>     7        6 X1=7, X2=6, X3=7 ~X1*X3 + X1*X2 0.931 0.818           1
#>     6        7 X1=6, X2=7, X3=7             X3 0.864 0.559           1
#>     7        7 X1=6, X2=7, X3=7             X3 0.864 0.576           1
#>     6        8 X1=7, X2=7, X3=7     X3 + X1*X2 0.906 0.853           1
#>     7        8 X1=7, X2=7, X3=7     X3 + X1*X2 0.906 0.879           1
4.5 Notes on output structure and reproducibility
All sweep functions return structured S3 objects containing:
  • Summary tables for cross-threshold comparison
  • Complete parameter settings
  • Detailed QCA objects (when return_details = TRUE)
In this paper, we show only summary outputs to maintain clarity and conciseness. Full details are nonetheless computed internally and remain accessible for reproducibility and further inspection.
5 Demonstration of advanced output and reporting capabilities
This section demonstrates the advanced output structures and reporting utilities of TSQCA. The purpose is technical: to illustrate how TSQCA organizes and reports results from threshold-sweep analyses. All demonstrations use the bundled dataset (sample_data) to ensure reproducibility.
library(TSQCA)
data("sample_data", package = "TSQCA")
5.1 Structured result objects
All TSQCA sweep functions return structured S3 objects. These objects store (i) analysis parameters (thresholds, cutoffs, and directional expectations) and (ii) cross-threshold summaries of solution structures and fit measures. When return_details = TRUE, they additionally retain detailed threshold-specific QCA objects (e.g., truth tables and solution objects) for inspection and reporting.
The following example illustrates the structure returned by otSweep().
res_ots <- otSweep(
  dat            = sample_data,
  outcome        = "Y",
  conditions     = c("X1", "X2", "X3"),
  sweep_range    = 6:8,
  thrX           = c(X1 = 7, X2 = 7, X3 = 7),
  include        = "?",
  dir.exp        = c(1, 1, 1),
  return_details = TRUE
summary(res_ots)
#> OTS-QCA Summary
#> Analysis Parameters:
#>   Outcome: Y
#>   Conditions: X1, X2, X3
#>   Consistency cutoff: 0.8
#>   Frequency cutoff: 1
#> Results by Threshold:
#>  thrY  expression inclS  covS n_solutions
#>     6  X3 + X1*X2 0.906 0.853           1
#>     7  X3 + X1*X2 0.906 0.879           1
#>     8 No solution    NA    NA           0
Although only summary information is printed here, the object retains the full set of threshold-specific results when requested (via return_details = TRUE). These include truth tables, minimized solutions, and associated fit measures.
5.2 Handling multiple solutions
At a given threshold, Boolean minimization may yield more than one logically equivalent solution. When this occurs, TSQCA preserves the multiplicity and internal structure of the solution set rather than collapsing it prematurely.
# Number of solutions identified at each threshold
res_ots\$summary[, c("thrY", "n_solutions")]
#>   thrY n_solutions
#> 1    6           1
#> 2    7           1
#> 3    8           0
In cases with multiple solutions, TSQCA distinguishes between EPIs, which are terms shared across all solutions, and SPIs, which vary across solutions. This distinction is maintained within the result object and can be reflected in downstream summaries and configuration tables.
5.3 Configuration chart generation
Beyond single-threshold summaries, TSQCA supports cross-threshold comparison of sufficient configurations. The primary representation is a sweep-level configuration chart, inspired by Fiss (2011), that aligns configurations across outcome thresholds: Columns correspond to outcome thresholds and rows correspond to conditions. Standardized symbols are used to encode the role of each condition within the sufficient configurations identified at each threshold: $\bullet$ indicates condition presence, $\otimes$ indicates condition absence, and a blank cell indicates “don’t care”.
This cross-threshold chart provides a compact visualization of how sufficient configurations emerge, persist, or disappear as the outcome threshold changes and thus constitutes a central output of TSQCA. Table 5 illustrates this format using the results from the otSweep() analysis presented earlier.
In the illustrative dataset used here, the identified configurations are relatively simple and involve only positive conditions, yielding a single solution at lower thresholds. This simplicity reflects properties of the data rather than limitations of the TSQCA framework. More complex patterns—including negated conditions and multiple equivalent solutions—are fully supported by the underlying output structures and reporting mechanisms.
Table 5: Cross-threshold configuration chart of sufficient conditions (solution-term level)
thrY = 6 (M1) thrY = 6 (M2) thrY = 7 (M1) thrY = 7 (M2)
X1 $\bullet$ $\bullet$
X2 $\bullet$ $\bullet$
X3 $\bullet$ $\bullet$
$\bullet$ = condition present; $\otimes$ = condition absent; blank = “don’t care”.
The symbol set can be switched to Unicode, ASCII, or LaTeX formats to accommodate different output environments.
Configuration charts are generated at the solution-term level by default, following Fiss (2011) notation, where each column represents one prime implicant. Threshold-level summary charts, which aggregate all configurations at each threshold into a single column, are also available.
5.4 Automated report generation
The generate_report() function produces structured Markdown reports directly from TSQCA result objects. These reports combine summary tables, configuration charts, and reproducibility information in a single document.
generate_report(
  result        = res_ots,
  output_file   = "tsqca_demo_report.md",
  title         = "TSQCA Demonstration Report",
  dat           = sample_data,
  format        = "full",
  include_chart = TRUE,
[…]
#> Report generated: tsqca_demo_report.md
The generated report can be used as:
  • A reproducible appendix
  • A starting point for further customization
5.5 Scope and limitations
This section is intentionally limited to demonstrating output and reporting capabilities. No substantive claims are made about the empirical meaning of the results derived from sample_data.
The primary objective is to show how TSQCA supports transparent, reproducible, and scalable reporting of threshold-sweep analyses within the standard R Markdown workflow.
6 Comparison with existing packages
6.1 Feature comparison
Table 6 compares TSQCA with existing QCA-related packages.
Table 6: Feature comparison ($\bullet$: Native implementation; $\circ$: Available via QCA package; —: Not available)
Feature TSQCA QCA SetMethods
Core QCA Functions
   Truth table construction $\circ$ $\bullet$ $\circ$
   Boolean minimization $\circ$ $\bullet$ $\circ$
   Necessity analysis $\circ$ $\bullet$ $\bullet$
Threshold Sweep
   Outcome threshold sweep $\bullet$
   Condition threshold sweep $\bullet$
   Dual threshold sweep $\bullet$
Output & Visualization
   Configuration table generation $\bullet$
   Markdown/LaTeX export $\bullet$
   EPI/SPI identification $\bullet$
6.2 Positioning
TSQCA is positioned as a complementary extension to the QCA ecosystem:
  • QCA package: Provides core computational infrastructure (truth tables, minimization)
  • SetMethods: Provides advanced QCA procedures (theory evaluation, MMR, visualization)
  • TSQCA: Provides systematic threshold exploration (sweep analysis, hierarchical sufficiency)
The three packages serve different but complementary purposes. Researchers can use QCA for standard analysis, SetMethods for robustness checks based on chosen thresholds, and TSQCA for exploring how results change across threshold ranges.
TSQCA is not intended as a robustness-testing alternative to SetMethods, but as an exploratory tool that is naturally used before or alongside confirmatory robustness checks.
Unique output capabilities: While QCA and SetMethods focus on analytical functions, TSQCA additionally provides publication-ready output generation. Configuration tables in Fiss notation ($\bullet$/$\otimes$) can be generated directly in LaTeX format for journal submissions, and comprehensive Markdown reports facilitate transparent documentation of threshold sensitivity analyses—features not available in existing packages.
6.3 Workflow comparison
Traditional workflow (without TSQCA):
# Manual iteration required
results <- list()
for (thr in 5:9) {
  # Manually dichotomize data
  dat\$Y_bin <- ifelse(dat\$Y >= thr, 1, 0)
  dat\$X1_bin <- ifelse(dat\$X1 >= 7, 1, 0)
  # ... repeat for all conditions
  tt <- QCA::truthTable(dat, outcome = "Y_bin", ...)
  sol <- QCA::minimize(tt)
  results[[as.character(thr)]] <- sol
  # Manual aggregation required
TSQCA workflow:
# Single function call
result <- otSweep(
  dat = dat,
  outcome = "Y",
  conditions = c("X1", "X2", "X3"),
  sweep_range = 5:9,
  thrX = c(X1 = 7, X2 = 7, X3 = 7)
# Structured output ready for analysis
7 Discussion
7.1 Limitations
[…]
Interpretation burden: More results require more interpretation. Researchers must develop frameworks for making sense of solution patterns across thresholds.
Data requirements: Meaningful threshold sweeps require sufficient variation in the data. If most cases cluster around a limited range of values, varying thresholds may not induce substantive changes in set membership or sufficient configurations.
7.2 Best practices
Based on our experience, we recommend the following practices:
  1. 1.
    Start with otSweep: Begin by exploring outcome threshold sensitivity before investigating condition thresholds.
  2. 2.
    Use meaningful threshold ranges: Select threshold ranges that correspond to substantively meaningful categories.
  3. 3.
    Document threshold rationale: Explicitly state why specific thresholds were chosen and what they represent substantively.
  4. 4.
    Report sensitivity results: Include threshold sensitivity analysis as a standard part of QCA reporting.
  5. 5.
    Interpret patterns, not just solutions: Focus on how solution structures evolve across thresholds, identifying recurring patterns such as hierarchical sufficiency.
8 Conclusion
8.1 Contributions
We introduce TSQCA, an R package that reconceptualizes thresholds in Qualitative Comparative Analysis as explicit objects of systematic exploration rather than fixed calibration choices. By enabling threshold sweep analyses through a unified and reproducible interface, TSQCA makes it possible to map how sufficient condition structures vary across substantively meaningful outcome and condition levels.
The TSQCA package makes three primary contributions:
  1. 1.
    Methodological innovation: We transform thresholds from fixed assumptions into objects of systematic study, enabling researchers to examine how causal structures emerge, disappear, and reorganize as thresholds vary.
  2. 2.
    Practical tools: Four sweep functions and automatic report generation reduce tedious manual processes to efficient single-function calls.
  3. 3.
    Analytical perspective: We demonstrate, through fully reproducible examples, how threshold variation can be examined systematically and summarized in a structured manner, without relying on substantive interpretation.
8.2 Conceptual implications for QCA practice
These implications are conceptual rather than formal theoretical propositions.
The hierarchical sufficiency pattern illustrated in this paper highlights a descriptive relationship between outcome stringency and causal complexity. Higher outcome levels typically requires more demanding combinations of conditions, whereas broader explanatory coverage is associated with more modest targets. This relationship provides a framework for understanding how causal complexity varies with outcome stringency.
The trade-off between coverage and consistency across thresholds offers researchers explicit guidance for target-setting decisions. Rather than arbitrarily selecting a single threshold, TSQCA enables informed choices based on a comprehensive mapping of the threshold-solution space.
8.3 Future directions
Several avenues for future development merit attention:
  1. 1.
    Computational efficiency: Reducing computational burden for large threshold spaces through parallel processing and algorithmic optimization.
  2. 2.
    Graphical visualization: Developing visualization functions such as heatmaps and trajectory plots to complement existing output capabilities.
  3. 3.
    Fuzzy-set extension: Adapting the threshold sweep framework for fsQCA, where calibration anchors (full membership, crossover point, full non-membership) can be systematically varied.
  4. 4.
    Integration with SetMethods: Combining TSQCA’s exploratory threshold analysis with SetMethods’ confirmatory robustness protocols to provide comprehensive robustness verification frameworks.
8.4 Concluding remarks
TSQCA represents a first step toward standardizing threshold sensitivity analysis in QCA research. By making such analyses accessible through user-friendly functions and structured outputs, we aim to elevate threshold examination from an occasional robustness check to a core component of QCA practice.
We encourage researchers to incorporate threshold sensitivity analysis as a standard element of QCA reporting. Doing so enhances the transparency, reliability, and credibility of set-theoretic causal inferences in social science research.
Acknowledgments
  • This work was supported by JSPS KAKENHI (Grant Number JP20K01998).
  • The author used Claude (Anthropic) for Japanese-to-English translation assistance and Editage for English language editing. The author takes full responsibility for all content.
9 Computational details
All analyses were performed using:
R version 4.5.2 (2025-10-31)
Platform: aarch64-apple-darwin20
Running under: macOS Tahoe 26.2
attached base packages:
[1] stats     graphics  grDevices utils     datasets  methods   base
other attached packages:
[1] QCA_3.23         admisc_0.39      TSQCA_1.2.0
[4] rmarkdown_2.30   rjtools_1.0.18.1
Execution times: The demonstrations in Section 4 completed in approximately 8 seconds total on the test system. Execution times scale with the number of threshold combinations: otSweep (3 thresholds, $\sim$0.8s), ctSweepS (4 thresholds, $\sim$1.1s), ctSweepM (8 combinations, $\sim$2s), and dtSweep (16 combinations, $\sim$4s). Times vary by hardware configuration.
Data and code availability:
Appendix: Function reference
This appendix provides a quick reference summary of all TSQCA functions. Table 7 lists each function with its purpose and key arguments for ease of lookup.
Table 7: TSQCA function summary
Function Purpose Key Arguments
otSweep() Outcome threshold sweep dat, outcome, conditions, sweep_range, thrX, dir.exp
ctSweepS() Single condition sweep dat, outcome, conditions, thrY, sweep_var, sweep_range, dir.exp
ctSweepM() Multi-condition sweep dat, outcome, conditions, thrY, sweep_list, dir.exp
dtSweep() Dual threshold sweep dat, outcome, conditions, sweep_range_Y, sweep_list_X, dir.exp
generate_report() Report generation result, output_file, title, dat